OpenStudy (anonymous):

HELP - THIS IS ON MY TEST TOMORROW: Factor Completely: 1. r^2t^2+10rt+25 2. 49m^2-14m+1 3. 4a^2-b^2

OpenStudy (anonymous):

for the first one, write it out as: \(\large (rt)^2+10rt + 25 \) so notice it looks like this: \(\large x^2+10x + 25 \) which can be factored: \(\large x^2+10x + 25 = (x+5)(x+5)=(x+5)^2\) so the original problem is actually: \(\large (rt + 5)^2 \)

OpenStudy (anonymous):

Give me a moment just reading over it now

OpenStudy (anonymous):

I'm confused as to how you wrote it. What is large?

OpenStudy (anonymous):

what's large? where?

OpenStudy (callisto):

I think that was a latex command :|

OpenStudy (anonymous):

oh... yeah...

OpenStudy (anonymous):

isn't the LaTex showing properly on your side rosimary?

OpenStudy (anonymous):

Can you retype it or use the equation button so that I can understand what you are explaining. PLEASE

OpenStudy (anonymous):